# A Class of Countably Covered Two-Dimensional Sofic Shifts

### v1

@@ Abstract
A multidimensional sofic shift is called countably covered if it has an SFT cover containing only countably many configurations. In contrast to the one-dimensional setting, not all countable sofic shifts are countably covered. We study a subclass of countable shift spaces and characterize the countably covered sofic shifts among them.
@@ 1 Introduction
A multidimensional shift of finite type (SFT) is a set of colorings of $\mathbb{Z}^{d}$ using finitely many colors that is defined by a finite number of local rules. A sofic shift is the image of an SFT, called its cover, under a map that identifies some of the colors with each other. We study the following problem: given a two-dimensional sofic shift whose cardinality is countable, when does it admit an SFT cover that is also countable?
@@
We concentrate on the case where $d=2$ and the sofic shift is vertically constant, that is, each coordinate of a vertical column must have the same color. This is the setting in which the author previously proved in [10] that not all countable sofic shifts have countable SFT covers. The author proved a necessary condition and showed it to be sufficient it in a restricted subclass of sofic shifts. In this article we prove the same statement in another subclass: those in which all but finitely many columns have the same “zero” color, and the number of nonzero columns is restricted by a computable function of the size of the smallest gap between two non-null columns.
@@
It is a well known result of Aubrun-Sablik [1] and Durand-Romashchenko-Shen [4] that the $\mathbb{Z}^{2}$-lift of a $\mathbb{Z}$-shift space is sofic, i.e. has an SFT cover, if and only if it is effectively closed. In [3], Durand and Romashchenko further prove that if the $\mathbb{Z}$-shift space is quasiperiodic, then the SFT cover can be taken quasiperiodic as well, and the same holds for minimality. One can say that quasiperiodicity and minimality are properties that can always be lifted to the SFT cover of a vertically constant sofic shift. Countability, as it turns out, cannot always be lifted in this way.
@@
On the other hand, this study is also related to the equal entropy problem of whether every $\mathbb{Z}^{2}$-sofic shift has an SFT cover of equal topological entropy. Desai showed in [2] that there always exist covers whose entropy is arbitrarily close to that of the sofic shift, but reaching the exact value remains an open question. In particular, it is not known whether every zero entropy sofic shift admits a zero entropy SFT cover, that is, whether having zero entropy is a liftable property. Coutability, like having zero entropy, is a “smallness property” of shift spaces, whose liftability we investigate.
@@
This article is dedicated to prof. Jarkko Kari, my former PhD advisor, on his 60th birthday. I am thankful to him for many years of collaboration and friendship.
@@ 2.1 Definitions
We recall some basic notions of multidimensional symbolic dynamics. Let $A$ be a finite alphabet and $d\geq 1$ a dimension. The $d$-dimensional full shift over $A$ is the set $A^{\mathbb{Z}^{d}}$, whose element are called configurations. It is equipped with the shift action of $\mathbb{Z}^{d}$, defined by $\sigma^{\vec{v}}(x)_{\vec{n}}=x_{\vec{n}+\vec{v}}$ for $x\in A^{\mathbb{Z}^{d}}$ and $\vec{v},\vec{n}\in\mathbb{Z}^{d}$. A $d$-dimensional pattern with finite domain $D\subset\mathbb{Z}^{d}$ is a map $P:D\to A$. Each pattern defines a cylinder set $[P]=\{x\in A^{\mathbb{Z}^{d}}\mid x|_{D}=P\}$, and these sets form a basis for the profinite topology. A topologically closed and shift-invariant subset of $A^{\mathbb{Z}^{d}}$ is called a shift space. If $X\subseteq Y$ are shift spaces, we say $X$ is a subshift of $Y$.
@@
A configuration $x\in A^{\mathbb{Z}^{2}}$ is vertically constant if $\sigma^{(0,1)}(x)=x$. A shift space is vertically constant if all its configurations are. The (two-dimensional) lift of a one-dimensional configuration $x\in A^{\mathbb{Z}}$ is the vertically constant configuration $x^{\dagger}\in A^{\mathbb{Z}^{2}}$ with $(x^{\dagger})_{(i,j)}=x_{i}$ for all $(i,j)\in\mathbb{Z}^{2}$. The lift of a shift space $X\subseteq A^{\mathbb{Z}}$ is $X^{\dagger}=\{x^{\dagger}\mid x\in X\}$.
@@
For $d\geq 2$, every effectively closed vertically constant $\mathbb{Z}^{d}$-shift space is sofic.
@@ 2.2 Previous work
In [10] it was shown that the converse holds if $X$ is a subshift of a one-dimensional countable sofic shift. Note that one-dimensional countable sofic shifts have a very constrained structure: their configurations consist of periodic regions of bounded periods, with a bounded number of period-breaking patterns in between [7, Lemma 4.8].
@@ 3 Gap width shifts
We define a class of countable one-dimensional shift spaces. For an alphabet $A$ with $0\in A$, denote $A_{0}=A\setminus\{0\}$. For a word $w\in A^{*}$, denote by $|w|_{\neq 0}$ the total number of nonzero symbols in $w$.
@@
Let $A$ be a finite alphabet containing 00, and $f:\mathbb{N}\to\mathbb{N}$ any function. The gap width shift $G(A,f)\subset\{0,1\}^{\mathbb{Z}}$ is the shift space defined by the forbidden patterns
$F(f)=\left\{w=ua0^{n}bv\in\{0,1\}^{*}\;\middle|\;n\geq 0,a,b\in A_{0},|w|_{\neq 0}>f(n)\right\}.$
If $A$ is clear from the context, we may denote $G(A,f)=G(f)$.
The gap width shift $G(f)$ contains exactly those configurations $x\in A^{\mathbb{Z}}$ for which the number of nonzero symbols is at most $f(n)$, where $n$ is the minimal length of a finite run of 0-symbols in $x$. In other words, the width of the smallest gap between nonzero symbols puts a bound on the number of gaps. It follows that $G(f)$ only contains configurations with a finite (but usually unbounded) number of nonzero symbols.
We are interested in subshifts of gap width shifts, and whether their lifts are sofic or countably covered. For the full gap width shifts $G(f)$, soficness of the lift corresponds to a computability-theoretic property of the function $f$.
@@
The lift $G(f)^{\dagger}$ is sofic if and only if $f:\mathbb{N}\to\mathbb{N}$ is upper semicomputable.
@@
By Theorem 1, $G(f)^{\dagger}$ is sofic if and only if $G(f)$ is effectively closed.
Suppose $f$ is upper semicomputable with computable approximation $g:\mathbb{N}^{2}\to\mathbb{N}$. Then $F(f)$ is computably enumerable, since the condition $|w|_{\neq 0}>f(n)$ is equivalent to $\exists k\in\mathbb{N}:|w|_{\neq 0}>g(n,k)$. Hence $G(f)$ is effectively closed.
Suppose then that $G(f)$ is effectively closed. Denote $F=\{0,1\}^{*}\setminus\mathcal{L}(G(f))$, which is a computably enumerable set. For each $n\in\mathbb{N}$ we have $f(n)=\min\left\{k\in\mathbb{N}\;\middle|\;a(0^{n}a)^{k}\in F\right\}$, where $a\in A_{0}$ is arbitrary. It follows that $f$ is upper semicomputable. ∎
As our main result, we show that the converse of Theorem 3 holds for subshifts of gap width shifts with computable gap functions. The proof is deferred to Section 5.
@@
Suppose then that $f$ is computable. Then $G(f)$ satisfies the countable cover conditions: it is effectively closed since the set of forbidden patterns $F(f)$ is computably enumerable, every configuration is eventually all-00 in both directions, and ${}^{\infty}uvw^{\infty}$ can be easily decided by measuring the smallest gap $110^{n}$ 1, computing $f(n)$ and comparing that to the number of $1$-symbols. ∎
In particular, the subclass of gap width shifts with countably covered sofic lifts is also characterized by a computability condition, and is neither empty nor equal to the class of gap width shifts with sofic lifts.
@@
Let $X\subseteq G(f)$ be a shift space. Define the gap function of $X$ by $f_{X}(n)=\max\{|x|_{\neq 0}\mid x\in X,m\leq n,a0^{m}b\sqsubset x\}$ for $n\geq 0$.
In other words, the gap function $f_{X}:\mathbb{N}\to\mathbb{N}$ is the pointwise smallest nondecreasing function such that $X\subseteq G(f_{X})$.
@@
As $X$ satisfies the countable cover conditions, we can algorithmically enumerate its configurations in the form ${}^{\infty}0v0^{\infty}$ for $v\in A^{*}$. Every such $v$ containing an occurrence of $a0^{n}b$ is a witness for $f_{X}(n)\geq|v|_{\neq 0}$, and $f_{X}(n)=|v|_{\neq 0}$ holds for at least one $v$. ∎
@@
There exists a subshift of a gap width shift that satisfies the countable cover conditions, but whose gap function is not computable. Theorem 6 does not apply in this situation; we do not know whether such shift spaces have countably covered lifts.
Define a shift space $X\subset\{0,1\}^{\mathbb{Z}}$ as the set of all translates of the following configurations: the all-00 configuration; configurations ${}^{\infty}0(10^{n})^{k}10^{\infty}$ for all $n\geq 2$ and $0\leq k\leq n$; and configurations ${}^{\infty}010^{n}(10^{k})^{k}10^{\infty}$ for all $n\geq 2$ such that the $n$th Turing machine halts on empty input in exactly $k>n$ steps. This set is a shift space. Since $f_{X}$ has Busy Beaver-like growth, it is not computable.

### v2

# On Countable SFT Covers of Sparse Multidimensional Shift Spaces

@@ Abstract
A multidimensional sofic shift is called countably covered if it has an SFT cover containing only countably many configurations. In contrast to the one-dimensional setting, not all countable sofic shifts are countably covered. We investigate the existence of countable covers for gap width shifts, where the number of nonzero symbols in a configuration is bounded by a function of the minimum distance between two such symbols. As our main results, we characterize those one-dimensional gap width shifts whose two-dimensional lift is a countably covered sofic shift, and show that a large class of two-dimensional gap width shifts are countably covered.
@@ 1 Introduction
A multidimensional shift of finite type (SFT) is a set of colorings of $\mathbb{Z}^{d}$ using finitely many colors that is defined by a finite number of local rules. A sofic shift is the image of an SFT, called its cover, under a map that identifies some of the colors with each other. We study the following problem: given a two-dimensional sofic shift whose cardinality is countable, when does it admit an SFT cover that is also countable? More generally, we may consider the problem of finding countable-to-one covers for arbitrary sofic shifts.
We concentrate on the case $d=2$ and the case where the sofic shift is in some sense “sparse”, that is, the alphabet contains 0 and each configuration contains few nonzero symbols: either their number is finite, or they lie on finitely many columns. In the latter case we also constrain each column to be uniform, so that each configuration is $(0,1)$-periodic and the number of configurations is necessarily countable.
@@
The existence of a global periodic direction is the setting in which the author previously proved in [12] that not all countable sofic shifts have countable SFT covers. The author proved a necessary condition and showed it to be sufficient it in a restricted subclass of sofic shifts. In this article we prove the same statement in another subclass: those in which all but finitely many columns have the same “zero” color, and the number of nonzero columns is restricted by a computable function of the size of the smallest gap between two nonzero columns. In particular, we can prove that not all sofic shifts of this form admit countable SFT covers.
@@
A two-dimensional shift space with $(0,1)$ as a global period is the lift of a one-dimensional shift space. It is a well known result of Aubrun-Sablik [1] and Durand-Romashchenko-Shen [5] that the $\mathbb{Z}^{2}$-lift of a $\mathbb{Z}$-shift space is sofic, i.e. has an SFT cover, if and only if it is effectively closed. In [4], Durand and Romashchenko further prove that if the $\mathbb{Z}$-shift space is quasiperiodic, then the SFT cover can be taken quasiperiodic as well, and the same holds for minimality. One can say that quasiperiodicity and minimality are properties that can always be lifted to the SFT cover of a vertically constant sofic shift. Countability, as it turns out, cannot always be lifted in this way.
@@
On the other hand, this study is also related to the equal entropy problem of whether every $\mathbb{Z}^{2}$-sofic shift has an SFT cover of equal topological entropy. Desai showed in [2] that there always exist covers whose entropy is arbitrarily close to that of the sofic shift, but reaching the exact value remains an open question. In particular, it is not known whether every zero entropy sofic shift admits a zero entropy SFT cover, that is, whether having zero entropy is a liftable property. Countability, like having zero entropy, is a “smallness property” of shift spaces, whose liftability we investigate.
Our second main result concerns truly sparse two-dimensional shift spaces, those for which the total number of nonzero symbols is bounded by a function of the minimum distance between two such symbols. We show that if there are no other constraints, and the function is upper semicomputable and bounded from above by $2^{\sqrt{n}}$, then the resulting shift space is a countably covered sofic shift. We do not provide examples of sparse two-dimensional sofic shifts that are not countably covered, as there are no known techniques for proving this property in the absence of a periodic direction.
These results are proved using a general toolbox of constructions for implementing computations and geometric or combinatorial constraints in countable shifts of finite type. We have presented them in a modular form for ease of reuse in other constructions of countable SFTs. Some of them may also be useful in constructing countable-to-one covers for general sofic shifts, although this is beyond the scope of this article.
The two main results, despite their seeming similarity, require subtly different approaches. The existence of a periodic direction (or rather, only having one non-periodic dimension) is both a blessing and a curse. On one hand, every finite pattern occurs in infinitely many positions, which allows us to ignore any given finite region of the configuration and check its contents somewhere else. On the other hand, every periodic configuration must have a periodic preimage in the SFT cover [12, Proposition 1], so we cannot enforce infinite computations in the cover. In both cases, the requirement of countability means that we cannot employ any of the known constructions of aperiodic SFTs, such as Robinson tiles [10] or fixed-point tiles [5], except in carefully controlled ways like restricting them to a finite region.
@@
This article is dedicated to prof. Jarkko Kari, my former PhD advisor, on his 60th birthday. I am thankful to him for many years of collaboration and friendship. I am also thankful to the anonymous referee for their comments that helped significantly improve this article.
@@ 2.1 Definitions
We recall some basic notions of multidimensional symbolic dynamics. Let $A$ be a finite alphabet containing 0 and $d\geq 1$ a dimension. Denote $A_{0}=A\setminus\{0\}$. The $d$-dimensional full shift over $A$ is the set $A^{\mathbb{Z}^{d}}$, whose element are called configurations. It is equipped with the shift action of $\mathbb{Z}^{d}$, defined by $\sigma^{\vec{v}}(x)_{\vec{n}}=x_{\vec{n}+\vec{v}}$ for $x\in A^{\mathbb{Z}^{d}}$ and $\vec{v},\vec{n}\in\mathbb{Z}^{d}$. A $d$-dimensional pattern with finite domain $D\subset\mathbb{Z}^{d}$ is a map $P:D\to A$. Each pattern defines a cylinder set $P]=\{x\in A^{\mathbb{Z}^{d}}\mid x|_{D}=P\}$, and these sets form a basis for the profinite topology. For a pattern $P$ over $A$, denote by $|P|_{\neq 0}$ the total number of nonzero symbols in $P$. A topologically closed and shift-invariant subset of $A^{\mathbb{Z}^{d}}$ is called a shift space. If $X\subseteq Y$ are shift spaces, we say $X$ is a subshift of $Y$.
@@
The language $\mathcal{L}(X)$ of a subshift $X\subseteq A^{\mathbb{Z}^{d}}$ is the set of finite patterns $P$ over $A$ such that $X\cap[P]\neq\emptyset$. If $X$ is effectively closed, then $\mathcal{L}(X)$ is a co-computably enumerable set.
A configuration $x\in A^{\mathbb{Z}^{2}}$ is vertically constant if $\sigma^{(0,1)}(x)=x$. A shift space is vertically constant if all its configurations are. The (two-dimensional) lift of a one-dimensional configuration $x\in A^{\mathbb{Z}}$ is the vertically constant configuration $x^{\dagger}\in A^{\mathbb{Z}^{2}}$ with $(x^{\dagger})_{(i,j)}=x_{i}$ for all $(i,j)\in\mathbb{Z}^{2}$. The lift of a shift space $X\subseteq A^{\mathbb{Z}}$ is $X^{\dagger}=\{x^{\dagger}\mid x\in X\}$. The lift $X^{\dagger}$ is effectively closed if and only if $X$ is.
@@
The two-dimensional lift $X^{\dagger}$ of a $\mathbb{Z}$-shift space $X$ is sofic if and only if $X$ is effectively closed.
@@ 2.2 Previous work
Note that these conditions imply that $X$ has a computable language. Namely, since $X$ is effectively closed, $\mathcal{L}(X)$ is co-computably enumerable. Conversely, if $P\in\mathcal{L}(X)$, then $P$ occurs in some $x\in X$ that is ultimately periodic in both directions, and by the third condition we can enumerate such configurations.
@@
In [12] it was shown that the converse holds if $X$ is a subshift of a one-dimensional countable sofic shift. Note that a one-dimensional countable sofic shift $Y\subseteq A^{\mathbb{Z}}$ has a very constrained structure: there is a number $p\geq 1$ such that every configuration $y\in Y$ satisfies $y_{i}=y_{i+p}$ for all $i\in\mathbb{Z}\setminus I_{y}$, where $I_{y}\subset\mathbb{Z}$ has cardinality at most $p$ [9, Lemma 4.8]. In other words, every configuration of $Y$ is locally periodic outside of a finite region (and the local periods and the size of the finite region, though not its diameter, are bounded).
@@ 3 Gap width shifts
3.1 Definition and main results
Recall that the max norm of a vector $\vec{v}=(v_{1},\ldots,v_{d})\in\mathbb{Z}^{d}$ is is $\|v\|_{\infty}=\max(|v_{1}|,\ldots,|v_{d}|)$, and the max distance of $\vec{u},\vec{v}\in\mathbb{Z}^{d}$ is $\|\vec{u}-\vec{v}\|_{\infty}$.
@@ 3.1 Definition and main results
Let $d\geq 1$, let $A$ be a finite alphabet containing 0, and let $f:\mathbb{N}\to\mathbb{N}$ be any function. The $d$-dimensional gap width shift $G_{d}(A,f)\subset A^{\mathbb{Z}^{d}}$ is the shift space defined by forbidding those $d$-dimensional patterns $P\in A^{D}$ such that for some $n\geq 0$ we have $|P|_{\neq 0}>f(n)$, and some $\vec{u}\neq\vec{v}\in D$ satisfy $P_{\vec{u}}\neq 0$, $P_{\vec{v}}\neq 0$ and $\|\vec{u}-\vec{v}\|_{\infty}\leq n$. We denote $G_{1}(A,f)=G(A,f)$. If $A$ is clear from the context, we may also denote $G_{d}(A,f)=G_{d}(f)$.
The gap width shift $G_{d}(f)$ contains exactly those configurations $x\in A^{\mathbb{Z}^{d}}$ for which the number of nonzero symbols is at most $f(n)$, where $n$ is the minimal max distance of two nonzero symbols in $x$. It follows that $G(f)$ only contains configurations with a finite (but usually unbounded) number of nonzero symbols.
We are interested in whether gap width shifts, their lifts or their subshifts are countably covered sofic shifts (or even sofic in the first place). For the one-dimensional full gap width shifts $G(f)$, soficness of the lift $G(f)^{\dagger}$ corresponds to a computability-theoretic property of the function $f$. It also gives a necessary condition of the soficness of multidimensional gap width shifts.
@@
Let $f:\mathbb{N}\to\mathbb{N}$ be arbitrary. For $d\geq 1$, the shift space $G_{d}(f)$ is effectively closed if and only if $f$ is upper semicomputable. The lift $G_{1}(f)^{\dagger}$ is sofic if and only if $f$ is upper semicomputable.
@@
Suppose $f$ is upper semicomputable with computable approximation $g:\mathbb{N}^{2}\to\mathbb{N}$. Then $F(f)$ is computably enumerable, since the condition $|P|_{\neq 0}>f(n)$ is equivalent to $\exists k\in\mathbb{N}:|P|_{\neq 0}>g(n,k)$. Hence $G(f)$ is effectively closed.
Suppose then that $G(f)$ is effectively closed. Let $F$ be the complement of $\mathcal{L}(G(f))$, which is a computably enumerable set. For each $n\in\mathbb{N}$, $f(n)$ is the smallest $k$ such that the $\{0,\ldots,kn\}\times\{0\}^{d-1}$-shaped pattern $a(0^{n-1}a)^{k}$ is in $F$, where $a\in A_{0}$ is arbitrary. It follows that $f$ is upper semicomputable.
The second claim follows from the first one and Theorem 1. ∎
As our first main result, we show that the converse of Theorem 3 holds for subshifts of one-dimensional gap width shifts with computable gap functions. The proof is deferred to Section 5.
@@
Suppose then that $f$ is computable. Then $G(f)$ satisfies the countable cover conditions: it is effectively closed since the set of forbidden patterns $F(f)$ is computably enumerable, every configuration is eventually all-0 in both directions, and $^{\infty}uvw^{\infty}$ can be easily decided by measuring the smallest gap $10^{n}1$, computing $f(n)$ and comparing that to the number of $1$-symbols. ∎
In particular, the subclass of one-dimensional gap width shifts with countably covered sofic lifts is also characterized by a computability condition, and is neither empty nor equal to the class of gap width shifts with sofic lifts.
Our second main result concerns two-dimensional gap width shifts. We show that a large subclass of them are in fact countably covered sofic shifts, including ones that satisfy the conditions of Proposition 13. The proof is given in Section 6.
Theorem 8.
Suppose that $f:\mathbb{N}\to\mathbb{N}$ is nondecreasing and upper semicomputable, and $f(n)<2^{\sqrt{n}}$ for all large enough $n\in\mathbb{N}$. Then $G_{2}(A,f)$ is a countably covered sofic shift.
Interestingly, the condition of the gap function is incomparable to that of Corollary 7. The computational restriction is weaker, intuitively because the lack of a periodic direction allows us to simulate arbitrary infinitely long computations. On the other hand, we are constrained by a fixed computable upper bound, which stems from the need to transmit information about the number of nonzero symbols in one region of a configuration to another. We know that the bound is suboptimal, and we have not seriously tried to optimize it since our efforts result in more complex constructions that still cannot reach the bound $2^{n}$.
3.2 Gap functions
The correspondence between the properties of subshifts of $G_{d}(f)$ and the computability-theoretic properties of the function $f$ is somewhat subtle. Let us investigate it in more detail.
@@ 3.2 Gap functions
Let $X\subseteq G_{d}(f)$ be a shift space. Define the gap function $f_{X}:\mathbb{N}\to\mathbb{N}$ of $X$ as follows: $f_{X}(n)$ is the smallest $k\geq 0$ such that there exists $x\in X$ with $|x|_{\neq 0}=k$ and $\min\{\|\vec{u}-\vec{v}\|_{\infty}\mid\vec{u}\neq\vec{v},x_{\vec{u}}\neq 0,x_{\vec{v}}\neq 0\}\geq n$.
In other words, the gap function $f_{X}:\mathbb{N}\to\mathbb{N}$ is the pointwise smallest nondecreasing function such that $X\subseteq G_{d}(f_{X})$. For simplicity, we concentrate on the one-dimensional case.
@@
As $X$ satisfies the countable cover conditions, we can algorithmically enumerate its configurations in the form $^{\infty}0v0^{\infty}$ for $v\in A^{*}$. Every such $v$ containing an occurrence of $a0^{n-1}b$ is a witness for $f_{X}(n)\geq|v|_{\neq 0}$, and $f_{X}(n)=|v|_{\neq 0}$ holds for at least one $v$. ∎
To recap, if $X=G(f)$ is effectively closed, then $f_{X}=f$ is upper semicomputable by Proposition 5. This does not hold for proper subshifts of $G(f)$: there exist effectively closed shift spaces $X\subseteq G(f)$ such that $f_{X}$ is not upper semicomputable. On the other hand, if $X\subseteq G(f)$ satisfies the countable cover conditions, then $f_{X}$ is lower semicomputable by Proposition 10. If we further know that $f_{X}$ is computable, then the lift $X^{\dagger}$ is a countably covered sofic shift by Theorem 6. But $f_{X}$ might not be computable (and a posteriori not upper semicomputable). The example below is a witness for both claims.
@@
There exists a subshift of a gap width shift that satisfies the countable cover conditions, but whose gap function is not bounded from above by any computable function. Theorem 6 does not apply in this situation; we do not know whether such shift spaces have countably covered lifts.
Define a shift space $X\subset\{0,1\}^{\mathbb{Z}}$ as the set of all translates of the following configurations: the all-0 configuration; configurations $^{\infty}0(10^{n})^{k}10^{\infty}$ for all $n\geq 2$ and $0\leq k\leq n$; and configurations $^{\infty}010^{n}(10^{k})^{k}10^{\infty}$ for all $n\geq 2$ such that the $n$th Turing machine halts on empty input in exactly $k>n$ steps. This set is a shift space. Since $f_{X}$ has Busy Beaver-like growth, it is not bounded from above by a computable function.
@@
3.3 The two-dimensional case
By Proposition 5, $f$ being upper semicomputable is a necessary condition for $G_{2}(A,f)$ being sofic. We do not know whether it is sufficient, although this seems plausible for the following reason. For a number $0\leq\alpha\leq 2$, let $X_{\alpha}\subset\{0,1\}^{\mathbb{Z}^{2}}$ be the shift space obtained by forbidding, for all $n\geq 2$, each $n\times n$ square with more than $n^{\alpha}$ occurrences of $1$. In his PhD thesis [3, Section III], Julien Destombes shows that $X_{\alpha}$ is sofic whenever $0\leq\alpha<1$ is an upper semicomputable number, and so is every effectively closed subshift of $X_{\alpha}$.
The construction uses fixed-point tilings introduced in [5], in which each configuration consists of a regular grid of $N_{1}\times N_{1}$ patterns called level-1 macrotiles, the macrotiles are arranged in a grid of $N_{1}N_{2}\times N_{1}N_{2}$ patterns called level-2 macrotiles, and so on. Each level-$k$ macrotile gathers information about the positions of $1$-symbols inside it, tries to find a forbidden pattern inside it, and if one was not found, passes said information to the level-$(k+1)$ macrotile containing it. The construction could presumably be modified so that each macrotile keeps track of just the number of $1$-symbols it contains and the smallest max distance between them, and checks whether the constraint of $G_{2}(A,f)$ is violated. Such a modification is beyond the scope of this article, as the fixed-point construction is very complex and never produces a countable-to-one cover.
On the other hand, even some seemingly tame two-dimensional gap width shifts have effectively closed subshifts that are not sofic. Hence, we cannot improve Theorem 8 to include arbitrary effectively closed subshifts of the $G_{2}(A,f)$ without further constraining the growth rate of $f$. To prove this, we need the following result.
Lemma 12 (Theorem 2.3 of [7], paraphrased).
For each sofic shift $X\subseteq A^{\mathbb{Z}^{2}}$ there exists $C>0$ with the following property. Let $m\geq 1$ and let $S\subseteq X$ have cardinality at least $C^{m}$. Then there exist $x\neq y\in S$ and a configuration $z\in X$ such that $z|_{[0,m-1]^{2}}=x|_{[0,m-1]^{2}}$ and $z|_{\mathbb{Z}^{2}\setminus[0,m-1]^{2}}=y|_{\mathbb{Z}^{2}\setminus[0,m-1]^{2}}$.
The simplest example that violates the conclusion of Lemma 12 is the mirror shift, a shift space over $\{0,1,2\}$ whose configurations contain a single infinite column of $2$-symbols surrounded by two half-planes over $\{0,1\}$ that are mirror images of each other (or no $2$-symbols). We implement a version of the mirror shift as a subshift of $G_{2}(\{0,1,2\},f)$.
Proposition 13.
Suppose that $f:\mathbb{N}\to\mathbb{N}$ is nondecreasing and computable, and that for all $C>0$ there exists $n\in\mathbb{N}$ with $f(n)>Cn$. Then there exists a computable subshift of $G_{2}(\{0,1,2\},f)$ that is not sofic.
Proof.
Given an integer $k\geq 1$, let $n_{k}\in\mathbb{N}$ be the smallest integer such that $f(k)\geq 1+2\lceil kn_{k}\log_{2}k\rceil$. The function $k\mapsto n_{k}$ is computable since $f$ is.
Define a subshift $X\subset G_{2}(\{0,1,2\},f)$ with the following constraints on a configuration $x\in X$. There can be at most one $2$-symbol. If $x$ contains at least two nonzero symbols with some distance $k\geq 1$, then within distance $kn_{k}$, at some position $\vec{v}$, there must be a $2$-symbol. If $x$ contains a $2$-symbol at position $\vec{v}$, then for some $k\geq 1$, every $1$-symbol is at a position $\vec{v}+(ik,jk)$ with $i\in[-n_{k},-1]\cup[1,n_{k}]$ and $j\in[0,n_{k}-1]$. Finally, there is a $1$-symbol at $\vec{v}+(ik,jk)$ if and only if there is one at $\vec{v}+(-ik,jk)$. This subshift is clearly computable.
We show that the conclusion of Lemma 12 does not hold for $X$. Let $C\geq 1$ be given. Let $T\subseteq\{(iC,jC)\mid i\in[-n_{C},-1]\cup[1,n_{C}],j\in[0,n_{C}-1]\}$ have cardinality $\lceil Cn_{C}\log_{2}C\rceil$. Let $R$ be the set of patterns of domain $0,Cn_{C}-1]^{2}$ that contain 0s or $1$s at positions of $T$ and 0s at all other positions. Each such pattern $P\in R$ can be extended into a configuration $x^{P}\in X$ in which $x^{P}_{(-C,0)}=2$ by placing a mirror image of $P$ to the left of the $2$-symbol and filling the rest of the configuration with 0-symbols. These extensions are in $G_{2}(\{0,1,2\},f)$, since the minimum distance between two nonzero symbols in $x^{P}$ is at least $C$, and $f(C)\geq 1+2\lceil Cn_{C}\log_{2}C\rceil$, which is the number of nonzero symbols in $x^{P}$.
Let $S=\{x^{P}\mid P\in R\}$. Then $|S|=|R|\geq 2^{\lceil Cn_{C}\log_{2}C\rceil}\geq C^{Cn_{C}}$, but there do not exist $x\neq y\in S$ satisfying the conclusion of Lemma 12, as all such $x$ and $y$ have different contents at $0,Cn_{C}-1]$, which are mirrored on the left of the $2$-symbol at (C,0)(-C,0). ∎